OpenStudy (itsonlycdeee):

A crew member on a fishing boat traveling due north off the coast of California observes that the bearing of Morro Rock from the boat is N 38° E. After sailing another 9.2 miles, the crew member looks back to find that the bearing of Morro Rock from the ship is S 23° E. At that time, how far is the boat from Morro Rock?

OpenStudy (itsonlycdeee):

 would this picture be correct in helping me find the answer?

OpenStudy (anonymous):

This is the picture I get. With this image you can calculate the other angle and the other 2 sides. Try that 

OpenStudy (itsonlycdeee):

or I'd use law of sines?

OpenStudy (anonymous):

I think you'd want to use the sin(38)=x/9.2 in this case because you want the distance from the second time (i.e. where the angle 23 is).

OpenStudy (itsonlycdeee):

@domu it says 5.7 is wrong :/ But there's another example where its N 86 degrees E and S 24 degrees E and the answer is 5.8.

OpenStudy (anonymous):

How did they generate the 5.8 answer there, cos or sin?

OpenStudy (anonymous):

Okie dokie, let's try the law of sines, what would be the set up then?

OpenStudy (itsonlycdeee):

SinA/a = SinB / b = SinC/c

OpenStudy (anonymous):

Yup, we don't need all 3 equalities, 2 will do. Now we just need to identify A,a, and B, where b is our unknown

OpenStudy (itsonlycdeee):

okay, so then A = 119 degrees & a = 9.2 mi, B = 23 degrees, b = unknown?